A firm with market power faces the following estimated demand and average variable cost functions:

Qd = 39,000 - 500P + 0.4M - 8,000PR

AVC = 30 - 0.005Q + 0.0000005Q2

where Qd is quantity demanded, P is price, M is income, and PR is the price of a related good. The firm expects income to be $40,000 and PR to be $2. Total fixed cost is $100,000. What is the estimated demand function for the firm?

Qd = 40,000 - 200P

Qd = 71,000 - 500P

Qd = 39,000 - 500P

Qd = 39,000 - 200P
Answer #1

Answer
Qd = 39,000 - 500P + 0.4M - 8,000PR

The firm expects income to be $40,000 and PR to be $2
Qd=39000-500P+0.4*40000-8000*2
Qd=39000-500P

the demand funtion is
Qd=39000-500P

Option 3